# Learning to Compress: Unlocking the Potential of Large Language Models for Text Representation

### v1

@@ Abstract
Text representation plays a critical role in tasks like clustering, retrieval, and other downstream applications. With the emergence of large language models (LLMs), there is increasing interest in harnessing their capabilities for this purpose. However, most of the LLMs are inherently causal and optimized for next-token prediction, making them suboptimal for producing holistic representations. To address this, recent studies introduced pretext tasks to adapt LLMs for text representation. Most of these tasks, however, rely on token-level prediction objectives, such as the masked next-token prediction (MNTP) used in LLM2Vec. In this work, we explore the untapped potential of context compression as a pretext task for unsupervised adaptation of LLMs. During compression pre-training, the model learns to generate compact memory tokens, which substitute the whole context for downstream sequence prediction. Experiments demonstrate that a well-designed compression objective can significantly enhance LLM-based text representations, outperforming models trained with token-level pretext tasks. Further improvements through contrastive learning produce a strong representation model (LLM2Comp) that outperforms contemporary LLM-based text encoders on a wide range of tasks while being more sample‑efficient, requiring significantly less training data.
@@ 1 Introduction
Text embedding models, which transform semantic content into dense vector representations, are fundamental to numerous tasks such as retrieval and recommendation. Early methods like TF-IDF and BM25 relied on simple statistics, thus falling short in capturing sequence semantics. The advent of deep neural networks led to the development of techniques such as Word2Vec (DBLP:journals/corr/abs-1301-3781). This paradigm shift paved the way for the foundational models like BERT (DBLP:conf/naacl/DevlinCLT19) and T5 (DBLP:journals/jmlr/RaffelSRLNMZLL20).
@@
Recently, there has been growing interest in leveraging the powerful capabilities of LLMs for text representation. However, most LLMs are inherently causal and optimized for next-token prediction, which makes them inherently suboptimal for generating holistic, coherent representations of entire sequences. To address this limitation, recent studies have proposed various pretext tasks for unsupervised adaptation of LLMs, focusing primarily on token-level prediction objectives. For example, LLM2Vec (DBLP:journals/corr/abs-2404-05961) first transforms the causal attention mechanism of LLMs into a bidirectional form, and then adopts masked next token prediction (MNTP) as a pretext task to align the training objectives of causal LLMs with those of bidirectional models such as BERT. In this setup, MNTP randomly masks tokens within a sentence and leverages contextual representations of preceding tokens to predict masked ones. On the other hand, Llama2Vec (DBLP:conf/acl/Li0XSL24) employs two pretext tasks, Embedding-Based Auto-Encoding (EBAE) and Embedding-Based Auto-Regression (EBAR), which predict tokens within the original sequence or the continued sequence. Although such objectives can capture the ”bag-of-tokens” information, they cannot fully preserve the coherent semantic integrity of the entire sequence. These tasks, therefore, remain fundamentally token-level rather than sequence-level prediction.
@@
In this work, we explore context compression as a pretext task for the unsupervised adaptation of LLMs to text encoders. Specifically, during compression pre-training, the model learns to produce compact “memory tokens” that replace the original context for downstream sequence prediction tasks. The conceptual difference between our objectives and previous pretext tasks is illustrated in Figure 1. Here, we explore several compression objectives: reconstruction task (DBLP:conf/iclr/00010WWCW24; DBLP:conf/acl/XuFMHLWZLTZZC24; DBLP:conf/nips/0002W00CWZ024) involves regenerating original sentences; continuation task (DBLP:conf/iclr/00010WWCW24; DBLP:conf/emnlp/ChevalierWAC23; DBLP:conf/acl/QinRCRD24; DBLP:conf/acl/XuFMHLWZLTZZC24; DBLP:journals/corr/abs-2401-07793) focuses on generating correct subsequent tokens. Our preliminary experiment shows that the reconstruction task does not provide satisfactory results, whereas the continuation task trained with NLL loss (CT-NLL) suffers from unstable training. Inspired by (DBLP:conf/nips/Mu0G23; DBLP:conf/emnlp/WingateSS22), we then propose Continuation Task with Knowledge Distillation (CTKD) as a pretext task, which aims to predict the probability of subsequent tokens in alignment with the original extended sequence from the soft prompt. Our experiments demonstrate that this well-designed objective (CTKD) can significantly enhance LLM-based text representations.
@@
  • We delve into the reasons behind the advantage of CTKD over other compression objectives, demonstrating that it suffers less from the dimensional collapse issue, making it more suited for downstream text representation.
@@ 2 Related Works
To better adapt LLMs for representation learning, a variety of training strategies have been proposed, including supervised contrastive learning (DBLP:journals/corr/abs-2402-05672; DBLP:conf/sigir/MaWYWL24; DBLP:journals/corr/abs-2402-09906; DBLP:conf/iclr/LiQXCLLSL25; DBLP:conf/acl/WangYHYMW24; DBLP:journals/corr/abs-2412-03223; DBLP:journals/corr/abs-2402-09906; DBLP:conf/acl/Li0XSL24; DBLP:journals/corr/abs-2404-05961), instruction tuning (DBLP:conf/acl/SuSKWHOYSZ023), and in-context learning (DBLP:conf/iclr/LiQXCLLSL25). Other research directions explore the use of synthetic data (DBLP:conf/acl/WangYHYMW24; DBLP:journals/corr/abs-2412-03223), multi-task learning (e.g., combining generation and representation learning) (DBLP:journals/corr/abs-2402-09906; DBLP:journals/corr/abs-2408-03402), or the use of pretext tasks (DBLP:conf/acl/Li0XSL24; DBLP:journals/corr/abs-2404-05961). Our work falls into the latter category, but focuses on the unique question of whether context compression can serve as an effective pretext task to adapt LLMs for text representation.
@@ 3.1 Method
$\hat{n}_{j}=g_{\phi}(\widetilde{m_{1}}\ldots,\widetilde{m_{k}},n_{1},\ldots,n_{j-1})$ (2)
@@
Biorxiv
Clustering
S2S
Medrxiv
Clustering
S2S
Twenty
newsgroups
Clustering
SciFact
NFCorpus
ArguAna
STS17
SICK-R
STS
Benchmark
Banking77
Classification
Emotion
Classification
Sprint
Duplicate
Questions
Stack
Overflow
DupQues.
SciDocs
$R$
75000 37500 59545 5483 3956 10080 5692 19854 2758 3696 2096 8931 82798 89131
@@ 3.2 Experiment Setup
The unsupervised adaptation of LLMs using the proposed compression-based pretext tasks is conducted on 32,000 samples from the English Wikipedia-103 dataset (DBLP:conf/iclr/MerityX0S17). This choice of dataset for unsupervised adaptation of LLMs to text embedders is consistent with the settings used in LLM2Vec and LLama2Vec. This dataset is chosen under the assumption that Wikipedia data is already included in the pre-training mixture of the LLM models we experiment with. Consequently, the adaptation process does not introduce new factual knowledge; rather, it focuses on teaching the model how to compress contextual information into soft tokens and construct sequence-level representations. For LLM2Comp, we select the default number of memory tokens to be 8, unless stated otherwise. We provide additional details of our training setup and hyperparameters in the Appendix  2.1.
@@ 4.2 Effective Token Index Before Meanpooling
To further verify the above hypothesis, we investigate how reducing correlated tokens impacts downstream performance. We define a token cluster as a set of tokens with high pairwise similarity. Initially, each token is treated as its cluster. Clusters are then merged if the minimum similarity between any pair of tokens across clusters exceeds 0.9. From each resulting cluster, we randomly select one representative token, and refer to these selected tokens as effective tokens. Sentence embeddings are then computed by applying mean pooling to the embeddings of these effective tokens. Using this procedure, the performance of LLM2Comp(NLL) trained on 128,000 samples increases from 48.38 to 51.09, as shown in Figure 7. We also include the performance of LLM2CompKL using a single compression token, as shown in Figure 7. The results indicate that the single-token setting performs unsatisfactorily, likely due to excessive information loss. This highlights the need for a more effective learning strategy that better balances information preservation and redundancy.
@@ 5.1 Method
Biorxiv
Clustering
S2S
Medrxiv
Clustering
S2S
Twenty
newsgroups
Clustering
SciFact
NFCorpus
ArguAna
STS17
SICK-R
STS
Benchmark
Banking77
Classification
Emotion
Classification
Sprint
Duplicate
Questions
Stack
Overflow
DupQues.
SciDocs
$R$
75000 37500 59545 5483 3956 10080 5692 19854 2758 3696 2096 8931 82798 89131
@@
$\textbf{Llm2Comp}_{KL}$ 0.36M Llama-2 36.53 32.85 53.14 75.05 39.13 58.20 91.61 83.05 85.33 82.99 52.24 96.16 51.45 84.85 65.90
@@
$\textbf{Llm2Comp}_{NLL}$ 0.36M Llama-2 37.15 33.70 55.11 76.79 39.72 59.37 91.40 83.32 86.31 84.57 54.01 96.27 51.90 85.36 66.78
@@ 5.4 Further Analysis
Figure 10: Performance across different training steps.
@@ 2.1 Training Details
Reranking (2)
StackOverflow
DupQuestions
82798

### v2

@@ Abstract
Text representation plays a critical role in tasks like clustering, retrieval, and other downstream applications. With the emergence of large language models (LLMs), there is increasing interest in harnessing their capabilities for this purpose. However, most of the LLMs are inherently causal and optimized for next-token prediction, making them suboptimal for producing holistic representations. To address this, recent studies introduced pretext tasks to adapt LLMs for text representation. Most of these tasks, however, rely on token-level prediction objectives, such as the masked next-token prediction (MNTP) used in LLM2Vec. In this work, we explore the untapped potential of context compression as a pretext task for unsupervised adaptation of LLMs. During compression pre-training, the model learns to generate compact memory tokens, which substitute the whole context for downstream sequence prediction. Experiments demonstrate that a well-designed compression objective can significantly enhance LLM-based text representations, outperforming models trained with token-level pretext tasks. Further improvements through contrastive learning produce a strong representation model (LLM2Comp) that outperforms contemporary LLM-based text encoders on a wide range of tasks while being more sample‑efficient, requiring significantly less training data. Code is available at https://github.com/longtaizi13579/LLM2Comp.
@@ 1 Introduction
Text embedding models, which transform semantic content into vector representations, are fundamental to numerous tasks such as retrieval and recommendation. Early methods like TF-IDF and BM25 relied on simple statistics, thus falling short in capturing sequence semantics. The advent of deep neural networks led to the development of techniques such as Word2Vec (DBLP:journals/corr/abs-1301-3781). This paradigm shift paved the way for the foundational models like BERT (DBLP:conf/naacl/DevlinCLT19) and T5 (DBLP:journals/jmlr/RaffelSRLNMZLL20).
@@
Recently, there has been growing interest in leveraging the powerful capabilities of LLMs for text representation. However, most LLMs are inherently causal and optimized for next-token prediction, which makes them inherently suboptimal for generating holistic, coherent representations of entire sequences. To address this limitation, recent studies have proposed various pretext tasks for unsupervised adaptation of LLMs, focusing primarily on token-level prediction objectives. For example, as shown in Figure 1(a), LLM2Vec (DBLP:journals/corr/abs-2404-05961) first transforms the causal attention mechanism of LLMs into a bidirectional form, and then adopts masked next token prediction (MNTP) as a pretext task to align the training objectives of causal LLMs with those of bidirectional models such as BERT. In this setup, MNTP randomly masks tokens within a sentence and leverages contextual representations of preceding tokens to predict masked ones. In contrast, as shown in Figure 1(b), Llama2Vec (DBLP:conf/acl/Li0XSL24) employs two pretext tasks, Embedding-Based Auto-Encoding (EBAE) and Embedding-Based Auto-Regression (EBAR), which predict tokens within the original sequence or the continued sequence. Although such objectives can capture the “bag-of-tokens” information, they cannot fully preserve the coherent semantic integrity of the entire sequence. These tasks, therefore, remain fundamentally token-level rather than sequence-level prediction.
@@
In this work, we explore context compression (Figure 1) as a pretext task for the unsupervised adaptation of LLMs to text encoders. Specifically, during compression pre-training, the model learns to produce compact “memory tokens” that replace the original context for downstream sequence prediction tasks. The conceptual difference between our objectives and previous pretext tasks is illustrated in Figure 1. Here, we explore several compression objectives: reconstruction task (DBLP:conf/iclr/00010WWCW24; DBLP:conf/acl/XuFMHLWZLTZZC24; DBLP:conf/nips/0002W00CWZ024) involves regenerating original sentences; continuation task (DBLP:conf/iclr/00010WWCW24; DBLP:conf/emnlp/ChevalierWAC23; DBLP:conf/acl/QinRCRD24; DBLP:conf/acl/XuFMHLWZLTZZC24; DBLP:journals/corr/abs-2401-07793) focuses on generating correct subsequent tokens. Our preliminary experiment shows that the reconstruction task does not provide satisfactory results, whereas the continuation task trained with NLL loss (CT-NLL) suffers from unstable training. Inspired by (DBLP:conf/nips/Mu0G23; DBLP:conf/emnlp/WingateSS22), we then propose Continuation Task with Knowledge Distillation (CTKD) as a pretext task, which aims to predict the probability of subsequent tokens in alignment with the original extended sequence from the soft prompt. Our experiments demonstrate that this well-designed objective (CTKD) can significantly enhance LLM-based text representations.
@@
  • We delve into the reasons behind the advantage of CTKD over other compression objectives, demonstrating that it is less prone to the dimensional collapse issue and thus more suited for downstream text representation.
@@ 2 Related Works
To better adapt LLMs for representation learning, a variety of training strategies have been proposed, including supervised contrastive learning (DBLP:journals/corr/abs-2402-05672; DBLP:conf/sigir/MaWYWL24; DBLP:journals/corr/abs-2402-09906; DBLP:conf/iclr/LiQXCLLSL25; DBLP:conf/acl/WangYHYMW24; DBLP:journals/corr/abs-2412-03223; DBLP:journals/corr/abs-2402-09906; DBLP:conf/acl/Li0XSL24; DBLP:journals/corr/abs-2404-05961; su-etal-2025-training), instruction tuning (DBLP:conf/acl/SuSKWHOYSZ023), and in-context learning (DBLP:conf/iclr/LiQXCLLSL25). Other research directions explore the use of synthetic data (DBLP:conf/acl/WangYHYMW24; DBLP:journals/corr/abs-2412-03223), multi-task learning (e.g., combining generation and representation learning) (DBLP:journals/corr/abs-2402-09906; DBLP:journals/corr/abs-2408-03402), variants of training loss functions (deng-etal-2025-following), or the use of pretext tasks (DBLP:conf/acl/Li0XSL24; DBLP:journals/corr/abs-2404-05961). Our work falls into the latter category, but focuses on the unique question of whether context compression can be utilized as an effective pretext task to enhance LLMs’ ability to learn text representations.
@@ 3.1 Method
$\hat{n}_{l}=g_{\phi}(\widetilde{m_{1}}\ldots,\widetilde{m_{k}},n_{1},\ldots,n_{l-1})$ (2)
@@
Bior. Medr. Twen. SciF. NFCo. Argu. STS17 SICK-R STSB. Bank. Emot. Spri. Stac. SciD.
@@ 3.2 Experiment Setup
The unsupervised adaptation of LLMs using the proposed compression-based pretext tasks is conducted on 32,000 samples from the English Wikipedia-103 dataset (DBLP:conf/iclr/MerityX0S17). This choice of dataset for unsupervised adaptation of LLMs to text embedders is consistent with the settings used in LLM2Vec and LLama2Vec. This dataset is chosen because Wikipedia is included in the pretraining data of the LLM models, thus the adaptation process does not introduce new factual knowledge; rather, it focuses on teaching the model how to compress contextual information into soft tokens and construct sequence-level representations. For LLM2Comp, we select the default number of memory tokens to be 8, unless stated otherwise. We provide additional details of our training setup and hyperparameters in the Appendix  2.1.
@@
  • MetaEOL (DBLP:conf/acl/LeiW00CTY24) designs eight meta-task prompts with ChatGPT-4 to guide LLMs to form sentence representations from multiple perspectives.
@@ 4.2 Effective Token Index Before Meanpooling
To further verify the above hypothesis, we investigate how reducing correlated tokens impacts downstream performance. We define a token cluster as a set of tokens with high pairwise similarity. Initially, each token is treated as its cluster. Clusters are then merged if the minimum similarity between any pair of tokens across clusters exceeds 0.9. From each resulting cluster, we randomly select one representative token, and refer to these selected tokens as effective tokens. Sentence embeddings are then computed by applying mean pooling to the embeddings of these effective tokens. Using this procedure, the performance of LLM2CompNLL trained on 128,000 samples increases from 48.38 to 51.09, as shown in Figure 7. We also include the performance of LLM2CompKL using a single compression token, as shown in Figure 7. The results indicate that the single-token setting performs unsatisfactorily, likely due to excessive information loss. This highlights the need for a more effective learning strategy that better balances information preservation and redundancy.
@@ 5.1 Method
Bior. Medr. Twen. SciF. NFCo. Argu. STS17 SICK-R STSB. Bank. Emot. Spri. Stac. SciD.
@@
$\textbf{LLM2Comp}_{KL}$ 0.36M Llama-2 37.15 33.70 55.11 76.79 39.72 59.37 91.40 83.32 86.31 84.57 54.01 96.27 51.90 85.36 66.78
@@
$\textbf{LLM2Comp}_{NLL}$ 0.36M Llama-2 36.53 32.85 53.14 75.05 39.13 58.20 91.61 83.05 85.33 82.99 52.24 96.16 51.45 84.85 65.90
@@ 5.3 Experimental Results
Figure 10: Performance across different training steps.
@@ 2.1 Training Details
Reranking (2) StackOverflow DupQuestions 82798